jhonyy9 (jhonyy9):

how can proveing that if a and b are natural numbers and p and k are primes so than for every a and b there are always one p and k such that a=(p-1)/2 and b=(k-1)/2 ?

OpenStudy (sriram):

tis not necessary simplifyin ur eq we get 2a=p-1 hence this wud be true for all cases where p is both odd n prime to give p-1=2a==even num it be valid if a=3 then p=7 true but if a=10 then p=21 not prime a=4 then p=9 not true either

jhonyy9 (jhonyy9):

so 1=(3-1)/2 2=(5-1)/2 ... ... n=(2n+1-1)/2 n=2n/2 n=n ... so hope so much that you like it ,... yes ?

jhonyy9 (jhonyy9):

... so but how can be proven that is true ?

jhonyy9 (jhonyy9):

... so i think by math induction

jhonyy9 (jhonyy9):

yes ? so if yes than ok but how please ?

OpenStudy (sriram):

it is NOT TRUE that every time p will be prime

OpenStudy (sriram):

thats what i m tryin to tell u u cant prove smth thats not true

jhonyy9 (jhonyy9):

why ? if p and k are always primes grater or equal 3

OpenStudy (sriram):

look at my first reply

OpenStudy (sriram):

for a =10 we dont get p as a prime number

OpenStudy (sriram):

for a =4 we dont get p as a prime no

jhonyy9 (jhonyy9):

3=2*1+1 --- 1=(3-1)/2 5=2*2+1 --- 2=(5-1)/2 7=2*3+1 --- 3=(7-1)/2 11=2*5+1 --- 5=(11-1)/2 13=2*6+1 --- 6=(12-1)/2 17=2*8+1 --- 8=(17-1)/2 19=2*9+1 --- 9=(19-1)/2

jhonyy9 (jhonyy9):

- ok - so than if this not is possibill getting for every natural numbers than from this result that for every natural numbers not exist one prim number for what we can writing that a=(p-1)/2 --- yes ? - so but we know that every primes can be writing in the form of 2n+1 - than ???